Professional baseball pitchers can deliver a pitch which can the blazing speed of 100 mph (miles per hour). A local team has reach drafted an up-and-coming mph. Assuming the ball has a mass 142.8 g and has this speed just before a batter would make contact, how much kinetic energy does the ball have? Number ball How high would the ball need to be dropped from to attain the same energy (neglect air resistance)? ohe acceleration due to gravity is 9.81 m/s2.) Number ball Previous ® Give ua s view solusan a Check Answer Next a Exit
Answer #1

We know that the kinetic energy = (1/2)mV2
where m is the mass of the ball , V is the velocity of the ball .
m = 142.8 gram = 0.1428 kg
V = 90 mph = 40.2336 m/s
Kinetic energy = (1/2)*(0.1428)*(40.2336)2 = 115.578 J
(2) We know that the potential energy = mgh
where m is the mass of the ball , h is height
mgh = 115.578
h = 82.505 m